OpenStudy (anonymous):

#ScienceQuestion2013

OpenStudy (♪chibiterasu):

#PutItInScience

OpenStudy (anonymous):

#NoThanks<3

OpenStudy (♪chibiterasu):

Plagiarism, everyone knows that :P

OpenStudy (anonymous):

mhm plagiarism ~(*o*)~

OpenStudy (♪chibiterasu):

Don't do dat shiz gurl ~(*o*)~

OpenStudy (anonymous):

got et

OpenStudy (anonymous):

.-.

OpenStudy (wolf1728):

Gee, since it seems we're never getting that science question, here's a science answer: E = mc²

OpenStudy (♪chibiterasu):

What about a couple of AlgeQuations: y = mx + b oooo and (y1 - y2) = m(x1 + x2)

OpenStudy (♪chibiterasu):

AlgeQuations, just made that up :3